OpenStudy (anonymous):

What is the range of the function y = x^2? all real numbers x ≥ 0 y ≥ 0

myininaya (myininaya):

what kinds of outputs do you get for x^2? like what is (-5)^2=? or (-1)^2=? or (0)^2=? or (5)^2=? can you describe what kinds of outputs we are getting?